OpenStudy (anonymous):

c) Calculate the Integral \( \Large \int_0^1 x^{3}dx \)

sam (.sam.):

Power rule

OpenStudy (anonymous):

can you implement it pls Sam ?

sam (.sam.):

\[\large \int\limits_{0}^{1}x^3dx\] You have to add 1 to the exponent of x, and then divide the exponent number which 'already' added 1 \[\large [\frac{x^{3+1}}{4}]_0^3\]

OpenStudy (anonymous):

sorry sam i am really bad at math, it would be much appreciate if you can do it for me.. I will give this up as a homework today..

OpenStudy (anonymous):

to my mentor

sam (.sam.):

where did you get confused?

OpenStudy (anonymous):

i understand what you meaned there but i dont know how to apply it..

OpenStudy (anonymous):

Sam i need to go now, thank you very much, if you write anything i will look when i am back see you later

sam (.sam.):

We'll do some examples \[\int\limits_{}^{}1~dx\] The technique here is to multiply 'x' into '1' you'll get =x then, you divide the whole thing (which is 'x') by the exponent of x \[=\frac{x}{1}\] then the final answer is =x ------------------------------------------------------ \[\int\limits_{}^{}x^2dx\] again, multiply 'x' into '\(x^2\)' \[=x(x^2)\] \[=x^3\] but we're not finished yet, you still need to divide the whole thing (which is '\(x^3\)') by the exponent of it, \[=\frac{x^3}{3}\] That's your final answer

OpenStudy (anonymous):

@.Sam. thank you very much..